OpenStudy (anonymous):

A trapezoid has an area of 166.5 in.^2 , height of 9 in. and one base 15 in. What is the length of the other base?

OpenStudy (michele_laino):

hint: the area A of a trapezoid, is given by the subsequent formula: \[A = \frac{{B + b}}{2} \times H\] where B is one base, b is the other base and H is the height

OpenStudy (michele_laino):

If I substitute your data into the above formula, I get: \[166.5 = \frac{{15 + x}}{2} \times 9\]

OpenStudy (michele_laino):

where x is the requested base

OpenStudy (michele_laino):

hint: next step is: \[15 + x = \frac{{166.5 \times 2}}{9} = ...?\]

OpenStudy (anonymous):

thank you, I think I got how to do it. Thank you so much

OpenStudy (michele_laino):

Thank you!